OpenStudy (anonymous):

Working at a constant rate, Bob can produce x/3 widgets in 8 minutes. Working at a constant rate, Jack can produce 2x widgets in 40 minutes, where x>0. Which of the following is greater: a. The number of minutes it will take Bob to produce 5x widgets. b. The number of minutes it will take Jack to produce 6x widgets.

OpenStudy (anikhalder):

Can you say how many minutes Jack will take to produce x widgets?

OpenStudy (anonymous):

8?

OpenStudy (anikhalder):

Time taken by Jack to produce 2x widgets = 40 minutes Hence, time taken by Jack to produce x widgets = ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

they're equal?

OpenStudy (anonymous):

2 minutes for both?

OpenStudy (anikhalder):

If you can make 20 objects in 40 minutes, how long will it take for you to make 1 object?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

2 inutes

OpenStudy (anonymous):

minutes

OpenStudy (anikhalder):

how did you calculate that?

OpenStudy (anonymous):

20/40 = 2/4 = 1/2 = 1 object every 2 minutes

OpenStudy (anikhalder):

okay so if Jack makes 2x objects in 40 minutes, how long will it take for him to make 1 object?

OpenStudy (anonymous):